Home Clean room MedAustron HEPHY testbeams old
testbeam-DESY-Oct2017 testbeam-DESY-April2017 testbeam-SPS2015 testbeam-SPS2014 testbeam-DESY14 testbeam-SPS12 testbeam-SPS11 testbeam-SPS10 testbeam-SPS09 testbeam-SPS08 LP-TPC
  HEPHY logbook for testbeam at SPS October 2011  Not logged in ELOG logo
Message ID: 41     Entry time: Thu Oct 6 15:57:51 2011
Author: Christian Irmler 
Project: BELLE 
Subject: Origami stack cooled 
Run Number: cal008 
Events: 83800 
StartTime: 06.10.2011 15:53:20 
EndTime: 06.10.2011 16:25:29  
Data: Good 

HV=+-40V,
CO2 cooling: P_TT_4: 18.2 bar, mf=1 g/s
beam on; xy table position x=306 y=76.5
I=19.5µA; old I2C settings; corrected trigger delay (htr); adjusted front-end voltages

Temp. at pipe inlet (K7) = -14°C
Temp. at pipe outlet (K8) = -17°C

Before this run, the front-end voltages (+2.5 and +1.25) were readjusted so that the nominal values appear on MAMBO. (Previously, both were about 50mV too low.) Not sure if this makes a big difference, let's see.

currents:
06.10.2011  15:54:24  event=2901  rate=36events/s temp/curr=35.5/-05.6 35.5/-06.0 35.4/-06.9 34.9/001.9 35.0/002.2 35.0/019.6
06.10.2011  16:24:25  event=81350  rate=50events/s temp/curr=34.9/-05.6 34.9/-06.1 34.9/-06.9 34.3/001.7 34.3/002.2 34.3/019.6
 

ELOG V3.1.4-966e3dd